OpenStudy (anonymous):

More complicated quadratic question.... Solve. Explain your steps. d^4+32=12d^2

OpenStudy (anonymous):

Your picture... :/

OpenStudy (anonymous):

This is the second time someone has commented on it in the last 10 minutes.....Its just a plant people lets be mature here.

OpenStudy (anonymous):

first lets summon all variables on the same side of equation d^4-12d^2+32=0

OpenStudy (jdoe0001):

have you covered quadratic equations yet?

OpenStudy (anonymous):

then to make it perfect square, we need to add 4 to 32 but not to distorb the original equation we need to subtract 4 as well

OpenStudy (anonymous):

I do not know how to solve quadratic equations what so ever.... Im horrible at them.

OpenStudy (anonymous):

d^4-12d^2+32+4-4=(d^4-12d^2+36)-4=0

OpenStudy (anonymous):

we can rewrite the paranthesis like (d^2-6)^2

OpenStudy (anonymous):

(d^2-6)^2-4=0 that is a perfect square as well

OpenStudy (anonymous):

so (d^2-6-2)(d^2-6+2)=0 to organize (d^2-8)(d^2-4)=0 either variables have to be equal to zero

OpenStudy (anonymous):

d^2-8=0 d=2\[\sqrt{2}\] d^2-4=0 d=2

OpenStudy (anonymous):

opps i mean \[2\sqrt{2}\]

OpenStudy (anonymous):

did you understand???